What is the osmotic pressure in atm of a 0.002m sucrose (C12H22011) solution at 20-degree celsius? use gas constant R=0.08210 L x atm/mol x k


1
Expert's answer
2021-12-17T05:44:02-0500

"\Pi=" icRT

Where "\Pi=" osmotic pressure

"i=" Van't Hoff index

"c=" molar concentration of solute

"R=" ideal gas constant

"T=" Temperature in Kelvin


"i=\alpha\>n+(1-\alpha)"


Sucrose does not dissociate in water

"\implies\>\alpha=0"

"\therefore\>i=1"


"\Pi=1×0.002×0.08210×(273+20)\\=0.048atm"